OpenStudy (anonymous):

Im taking 12th grade calculus and im only in 9th grade i dont know the answer to this question ?d^2y/dx^2=

OpenStudy (anonymous):

itis incomplete que.

OpenStudy (anonymous):

it is rather ugly notation, but it means take the second derivative that is take the derivative, then take the derivative of the result

OpenStudy (anonymous):

Take derivative twice!

OpenStudy (anonymous):

for example \[f(x)=x^2\] \[f'(x)=2x\] \[f''(x)=2\]

OpenStudy (anonymous):

\[f(x)=10x-\frac{4}{x}\] \[f'(x)=10+\frac{4}{x^2}\] \[f''(x)=-\frac{8}{x^3}\]

OpenStudy (anonymous):

what are you doing taking calc in 9th grade?! you must be some student.